# Context-Alignment: Activating and Enhancing LLM Capabilities in Time Series

### v1

@@ Abstract
Recently, leveraging pre-trained Large Language Models (LLMs) for time series (TS) tasks has gained increasing attention, which involves activating and enhancing LLMs’ capabilities. Many methods aim to activate LLMs’ capabilities based on token-level alignment, but overlook LLMs’ inherent strength on natural language processing — their deep understanding of linguistic logic and structure rather than superficial embedding processing. We propose Context-Alignment, a new paradigm that aligns TS with a linguistic component in the language environments familiar to LLMs to enable LLMs to contextualize and comprehend TS data, thereby activating their capabilities. Specifically, such context-level alignment comprises structural alignment and logical alignment, which is achieved by a Dual-Scale Context-Alignment GNNs (DSCA-GNNs) applied to TS-language multimodal inputs. Structural alignment utilizes dual-scale nodes to describe hierarchical structure in TS-language, enabling LLMs treat long TS data as a whole linguistic component while preserving intrinsic token features. Logical alignment uses directed edges to guide logical relationships, ensuring coherence in the contextual semantics. Demonstration examples prompt are employed to construct Demonstration Examples based Context-Alignment (DECA) following DSCA-GNNs framework. DECA can be flexibly and repeatedly integrated into various layers of pre-trained LLMs to improve awareness of logic and structure, thereby enhancing performance. Extensive experiments show the effectiveness of DECA and the importance of Context-Alignment across tasks, particularly in few-shot and zero-shot forecasting, confirming that Context-Alignment provide powerful prior knowledge on context. We will release the source code upon publication.
@@ 1 Introduction
Regarding the first issue, existing works are dedicated to align TS token embeddings with language token embeddings Jin et al. (2024); Sun et al. (2024); Pan et al. (2024). However, whether such token-level alignment can fully leverage the LLMs’ potential remains questionable. Inspired by recent researches on LLMs Ethayarajh (2019); Nie et al. (2024); Wang et al. (2023), we reconsider the inherent advantages of LLMs in natural language processing (NLP). We believe the strength of LLMs primarily stems from their deep comprehension of language logic and structure, rather than superficial token embedding processing. Clearly, the excessive accumulation of tokens without logical guidance often struggles to effectively convey meaning. Especially TS-language multimodal inputs are lengthy, lack structure and coherent semantics, greatly challenging LLMs’ comprehension. Regarding the second issue, current methods aim to directly enhance LLMs’ capabilities in TS tasks through techniques such as TS decomposition Cao et al. (2024) and optimizing prompts Chuang et al. (2024). However, without adequately addressing the first issue, these methods need more interpretability and the improvements remain limited. A natural solution to these issues is to fully leverage the strengths of LLMs to transform TS tasks into NLP-like tasks, activating LLMs capabilities firstly. Then, leveraging NLP techniques further enhances LLMs performance on TS tasks.
@@
In this paper, we propose Context-Alignment, a new paradigm that aligns TS data with a linguistic component in the language environment familiar with LLMs. Such context-level alignment leverages the LLMs’ inherent strength in logic and structure to enable LLMs to contextualize and comprehend TS data, thereby activating their capabilities. Context-Alignment contains structural alignment and logical alignment to construct a consistent context for TS-language multimodal inputs. We develop a Dual-Scale Context-Alignment Graph Neural Networks (DSCA-GNNs) framework to achieve both structural and logical alignment. Specifically, structural alignment employs dual-scale nodes to describe hierarchical structure in TS-language, i.e. the structural independence of tokens and the overall structure of modalities. Structural alignment provides LLMs with structural segmentation information for lengthy TS language inputs, enabling LLMs treat long TS data as a individual linguistic component while preserving intrinsic token features. Logical alignment uses directed edges in both scale GNNs to guide the local and global logical relationship between TS data and language prompts, integrating TS within the language environment and ensuring semantic coherence across two modalities. Utilizing the demonstration examples prompt technique Brown (2020), we propose Demonstration Examples based Context-Alignment (DECA) following the DSCA-GNNs framework, which further enhances the LLMs’ performance on TS tasks. DECA can be flexibly and repeatedly integrated into various layers of pre-trained LLMs to improve awareness of logic and structure. Extensive experiments across various TS tasks demonstrate the effectiveness of our method. Notably, in few-shot and zero-shot forecasting tasks, our approach significantly outperforms others, confirming that the logical and structural alignment provide powerful prior knowledge on context. Ablation studies further validate the importance of Context-Alignment.
In summary, our core contributions in this work can be summarized as below:
@@
  • We develop a Dual-Scale Context-Alignment GNNs framework, which achieves structural and logical alignment through dual-scale nodes and directed edges, thus realizing Context-Alignment. Furthermore, by integrating the Demonstration Examples Prompt technique, we introduce DECA, which enhances LLMs’ performance in TS tasks.
@@
Figure 1: The architecture of our method, where the graph structure demonstrates the prediction task based on DECA as detailed in Sec.3.3. Dual-Scale Context-Alignment GNNs can be flexibly and repeatedly integrated into pre-trained LLMs at various layers, enhancing LLMs’ awareness of logic and structure and improving performance.
@@ 2.1 Time Series Tasks
TS tasks are crucial for various applications, including financial forecasting, weather prediction, and activity recognition, and involve analyzing time-ordered data. Traditional statistical methods like ARIMA Anderson (1976) and Prophet Taylor and Letham (2018) effectively model trends and seasonality. The rise of deep learning introduced CNN-based methods, which use convolutional neural networks to extract features automatically Bai et al. (2018); Wu et al. (2022). Additionally, RNNs, such as LSTM and GRU, excel in sequence prediction by capturing dynamic temporal behaviors and long-range dependencies Lai et al. (2018a); Qin et al. (2017); Siami-Namini et al. (2018). More recently, Transformer-based models have advanced the field by processing sequences in parallel and applying attention mechanisms to focus on significant temporal aspects, thus improving performance in complex scenarios Wen et al. (2023); Zhou et al. (2022); Wu et al. (2021). However, these methods often struggle with long, intricate sequences and lack the flexibility and generalizability required for diverse real-world TS data across diverse domains.
@@ 2.2 Large Language Models for Time Series
LLMs have demonstrated strong capabilities in various fields, becoming a focal point for advancing TS tasks. Recent research highlights their potential in TS tasks. Some methods aim to enhance the capabilities of LLMs for TS tasks directly. For example, Cao et al. (2024) captures complex interactions among trend, seasonal, and residual components to aid distribution adaptation. Chuang et al. (2024) propose a statistical prompting strategy to enhance the performance. However, these methods overlook an important step: enable LLMs to understand TS inputs firstly. Other methods aim to enable LLMs to understand TS data, and activate their potential for TS tasks. For example,  Jin et al. (2024) reprograms input TS into text prototypes and enhances them with prompts. Sun et al. (2024) aligns TS embeddings space with LLMs embeddings space. Pan et al. (2024) proposes $S^{2}$-LLM, a semantic space-informed prompt learning to align TS embeddings with language token embeddings. However, these methods overlook that LLMs’ inherent strength stems from their understanding of the logic and structure, rather than superficial token embedding processing. Merely aligning token embeddings without considering the coherence and consistency of the context fails to leverage this inherent advantage.
@@ 3.2 Demo: Vanilla Context-Alignment (VCA)
The most straightforward approach to utilizing pre-trained LLMs for TS tasks is to input both TS data and vanilla language prompts into the model directly. For TS forecasting tasks, a vanilla prompt like “Predict future sequences using previous data:" can be employed to guide LLMs completing the prediction. The embeddings of the prompt can be represented as $\{{\bm{z}}_{1},{\bm{z}}_{2},\dots,{\bm{z}}_{m}\}$, and the overall input embeddings of the model can be represented as below:
@@
Due to the verbose and lacking in clear structural divisions of TS embeddings $\{\bm{e}_{i}\}_{i=1}^{n}$, LLMs lack information that TS embedding $\{\bm{e}_{i}\}_{i=1}^{n}$ is an integral entity, making it difficult to analyze TS. Furthermore, form 1 lacks logical guidance. Directly concatenating the TS embeddings and language prompt embeddings loses essential contextual coherence between TS data and the prompt.
Dual-Scale Context-Alignment GNNs of VCA. Context-Alignment leverages the LLMs’ inherent strength in logic and structure to enable LLMs to contextualize and comprehend TS data. We delineate clear structures (structural alignment) and establish correct logical guidance (logical alignment) by a Dual-Scale Context-Alignment GNNs framework to achieve Context-Alignment. Firstly, using dual-scale nodes, structural alignment aggregates tokens from the same modality into one linguistic component while preserving the feature of each token. Fine-grained GNN $G_{F}$ treats each token, i.e., each element in form 1 as a node. Coarse-grained GNN $G_{C}$ treats consecutive tokens with same modality as a node. Specifically, $G_{C}$ employs two learnable linear layers, $f_{e}$ and $f_{z}$, to embed the TS tokens and language tokens into an $M$-dimensional space, respectively, which can be formalized as
@@
Each element in form 2 is regarded as a node in $G_{C}$. Secondly, using directed edges, logical alignment emphasizes the correct semantic association between different components. Clearly, in the prompt “Predict future sequences using previous data:", “previous data:” refers to the lengthy TS data, and requires information from TS data. Therefore, in $G_{C}$, we construct directed edge $E:\tilde{\bm{e}}\to\tilde{\bm{z}}$ to indicate that the entire TS data serves as the upstream information source for the prompt. In $G_{F}$, we construct directed edges $\{E_{ij}:{\bm{e}}_{i}\to{\bm{z}}_{j}|i\in\{1,\dots,n\}$, $j\in\{1,\dots,m\}\}$ to convey varying information from each TS token to the prompt token. We also constrain the sum of edge weights from TS tokens to one language token to be $1$, i.e. $\sum_{i=1}^{n}w_{ij}=1$, $w_{ij}$ is the edge weight of $E_{ij}$, implicitly emphasizing that TS should be treated as a whole. In $G_{C}$, the weights of all directed edges are set to $1$. Besides, $\{w_{ij}\}_{i=1}^{n}$ are proportional to the cosine similarity between the embeddings of two nodes.
@@
Overview. After structural and logical alignment, $\hat{\bm{N}}_{F}$ maintains a clear structural and coherent semantics information, helping LLMs to comprehend TS tasks and activating potential capabilities. Both $\hat{\bm{N}}_{F}$ and $\hat{\bm{N}}_{C}$ are input into pre-trained LLMs. The DSCA-GNNs can be flexibly integrated into various layers of pre-trained LLMs, and only the first time apply it need $f_{e}$ and $f_{z}$ to obtain coarse-grained GNN. The output from $G_{F}$ branch is used to compute the MSE loss against the ground truth. In this section, we use the vanilla prompt to construct specific DSCA-GNNs framework, we call it Vanilla Context-Alignment (VCA).
3.3 Demonstration Examples Based Context-Alignment (DECA)
@@
In NLP, “demonstration examples prompt" refers to a small set of instances provided to the model to demonstrate a task, enabling the model to perform similar tasks effectively Brown (2020). Based on this, we divide the TS embeddings $\{\bm{e}_{i}\}_{i=1}^{n}$ into $N$ parts while preserving their original order. Since the subsequent TS part can be used as ground truth for predictions based on the preceding TS parts, we can construct $N-1$ prediction demonstration examples using $N$ parts of $\{\bm{e}_{i}\}_{i=1}^{n}$ and language prompt embeddings $\{{\bm{z}}_{1},{\bm{z}}_{2},\dots,{\bm{z}}_{m}\}$ same as demo in Sec. 3.2. The $j$-th part of $\{\bm{e}_{i}\}_{i=1}^{n}$ is denoted as $\{{\bm{e}}_{j,1},\dots,{\bm{e}}_{j,l_{j}}\}$. We arrange the TS-language embeddings in the format as:
@@
Dual-Scale Context-Alignment GNNs of DECA.  Due to the verbose and lacking in clear structural divisions of format 5, LLMs struggle to understand the input. We first construct the coarse-grained GNN $G_{C}$. We introduce two learnable linear layers $f_{e}$ and $f_{z}$ to embed the TS tokens and language tokens into an $M$-dimensional space, respectively, which can be formalized as
@@
Fine-grained GNN $G_{F}$ treats each token in form 5 as a node. The directed edge set of the $G_{F}$ can be represented as follows:
@@
Formula LABEL:edgeF means that the directed edges of $G_{F}$ is a decomposition of the directed edges in $G_{C}$. Additionally, since LLMs have strong comprehension abilities for language prompts, we can prune the directed edges in $G_{F}$, transforming form LABEL:edgeF into form LABEL:edgeF2, i.e., TS tokens are only connected to the first and last tokens of the prompt, thereby prevent overfitting. We constrain $\sum_{s=1}^{l_{j}}w^{(i)}_{j,s}=1$ for the first type edges, $\sum_{s=1}^{l_{i+1}}w^{(i)}_{i+1,s}=1$ for the second type edges, where $w^{(i)}_{j,s}$ is the edge weight of ${\bm{e}}_{j,s}\to{\bm{z}}^{(i)}_{1}$, $w^{(i)}_{i+1,s}$ is the edge weight of ${\bm{z}}^{(i)}_{m}\to{\bm{e}}_{i+1,s}$. $\{w^{(i)}_{i+1,s}\}_{s=1}^{l_{j}}$ and $\{w^{(i)}_{i+1,s}\}_{s=1}^{l_{i+1}}$ are proportional to the cosine similarity between node embeddings. Fig. 1 provides a schematic of the structure. The node embedding update formula is similar to formula 3.
@@
Learnable interactions & Overview. Similar to the demo in Sec. 3.2, we introduce an assignment matrix and a learnable weight matrix to achieve learnable interactions between the two scales. The training of DECA can refer to Sec. 3.2 and Fig. 1.

### v2

@@ Abstract
Recently, leveraging pre-trained Large Language Models (LLMs) for time series (TS) tasks has gained increasing attention, which involves activating and enhancing LLMs’ capabilities. Many methods aim to activate LLMs’ capabilities based on token-level alignment, but overlook LLMs’ inherent strength in natural language processing — their deep understanding of linguistic logic and structure rather than superficial embedding processing. We propose Context-Alignment (CA), a new paradigm that aligns TS with a linguistic component in the language environments familiar to LLMs to enable LLMs to contextualize and comprehend TS data, thereby activating their capabilities. Specifically, such context-level alignment comprises structural alignment and logical alignment, which is achieved by Dual-Scale Context-Alignment GNNs (DSCA-GNNs) applied to TS-language multimodal inputs. Structural alignment utilizes dual-scale nodes to describe hierarchical structure in TS-language, enabling LLMs to treat long TS data as a whole linguistic component while preserving intrinsic token features. Logical alignment uses directed edges to guide logical relationships, ensuring coherence in the contextual semantics. Following the DSCA-GNNs framework, we propose an instantiation method of CA, termed Few-Shot prompting Context-Alignment (FSCA), to enhance the capabilities of pre-trained LLMs in handling TS tasks. FSCA can be flexibly and repeatedly integrated into various layers of pre-trained LLMs to improve awareness of logic and structure, thereby enhancing performance. Extensive experiments show the effectiveness of FSCA and the importance of Context-Alignment across tasks, particularly in few-shot and zero-shot forecasting, confirming that Context-Alignment provides powerful prior knowledge on context. The code is open-sourced at https://github.com/tokaka22/ICLR25-FSCA.
@@ 1 Introduction
Regarding the first issue, existing works primarily focus on aligning TS token embeddings with language token embeddings (Jin et al., 2024; Sun et al., 2024; Pan et al., 2024). However, whether such token-level alignment can fully leverage the LLMs’ potential remains questionable. Inspired by recent research on LLMs (Ethayarajh, 2019; Nie et al., 2024; Wang et al., 2023), we reconsider the inherent advantages of LLMs in natural language processing (NLP). We believe the strength of LLMs primarily stems from their deep comprehension of language logic and structure, rather than superficial token embedding processing. Clearly, the excessive accumulation of tokens without logical guidance often struggles to effectively convey meaning. Especially TS-language multimodal inputs are lengthy, and lack structure and coherent semantics, greatly challenging LLMs’ comprehension. Regarding the second issue, current methods aim to directly enhance LLMs’ capabilities in TS tasks through techniques such as TS decomposition (Cao et al., 2024) and optimizing prompts (Chuang et al., 2024). However, without adequately addressing the first issue, these methods need more interpretability and the improvements remain limited. A natural solution to these issues is to fully leverage the strengths of LLMs to transform TS tasks into NLP-like tasks, activating LLMs’ capabilities first. Then, leveraging NLP techniques further enhances LLMs’ performance on TS tasks.
@@
In this paper, we propose Context-Alignment, a new paradigm that aligns TS data with a linguistic component in the language environment familiar with LLMs. Such context-level alignment leverages the LLMs’ inherent strength in logic and structure to enable LLMs to contextualize and comprehend TS data, thereby activating their capabilities. Context-Alignment contains structural alignment and logical alignment to construct a consistent context for TS-language multimodal inputs. We develop a Dual-Scale Context-Alignment Graph Neural Networks (DSCA-GNNs) framework to achieve both structural and logical alignment. Specifically, structural alignment employs dual-scale nodes to describe hierarchical structure in TS-language, i.e. the structural independence of tokens and the overall structure of modalities. Structural alignment provides LLMs with structural segmentation information for lengthy TS language inputs, enabling LLMs to treat long TS data as an individual linguistic component while preserving intrinsic token features. Logical alignment uses directed edges in both scale GNNs to guide the local and global logical relationship between TS data and language prompts, integrating TS within the language environment and ensuring semantic coherence across two modalities. Utilizing the few-shot prompting technique (Brown, 2020), we propose Few-Shot prompting based Context-Alignment (FSCA) following the DSCA-GNNs framework, which further enhances the LLMs’ performance on TS tasks. FSCA can be flexibly and repeatedly integrated into various layers of pre-trained LLMs to improve awareness of logic and structure. Extensive experiments across various TS tasks demonstrate the effectiveness of our method. Notably, in few-shot and zero-shot forecasting tasks, our approach significantly outperforms others, confirming that the logical and structural alignment provides powerful prior knowledge on context. Ablation studies further validate the importance of Context-Alignment.
In summary, our core contributions can be summarized below:
@@
  • We develop a Dual-Scale Context-Alignment GNNs framework, which achieves structural and logical alignment through dual-scale nodes and directed edges, thus realizing Context-Alignment. Furthermore, by integrating the few-shot prompting technique, we introduce (FSCA), which enhances LLMs’ performance in TS tasks.
@@
Figure 1: The architecture of our method, where the graph structure demonstrates the prediction task based on FSCA as detailed in Sec.3.3. Dual-Scale Context-Alignment GNNs can be flexibly and repeatedly integrated into pre-trained LLMs at various layers, enhancing LLMs’ awareness of logic and structure and improving performance.
@@ 2.1 Time Series Tasks
TS tasks are crucial for various applications, including financial forecasting, weather prediction, and activity recognition, and involve analyzing time-ordered data. Traditional statistical methods like ARIMA (Anderson, 1976) and Prophet (Taylor and Letham, 2018) effectively model trends and seasonality. The rise of deep learning introduced CNN-based methods, which use convolutional neural networks to extract features automatically (Bai et al., 2018; Wu et al., 2022). Additionally, RNNs, such as LSTM and GRU, excel in sequence prediction by capturing dynamic temporal behaviors and long-range dependencies (Lai et al., 2018a; Qin et al., 2017; Siami-Namini et al., 2018). More recently, Transformer-based models have advanced the field by processing sequences in parallel and applying attention mechanisms to focus on significant temporal aspects, thus improving performance in complex scenarios (Wen et al., 2023; Zhou et al., 2022; Wu et al., 2021). However, these methods often struggle with long, intricate sequences and lack the flexibility and generalizability required for diverse real-world TS data across diverse domains.
@@ 2.2 Large Language Models for Time Series
LLMs have demonstrated strong capabilities in various fields, becoming a focal point for advancing TS tasks. Recent research highlights their potential in TS tasks. Some methods aim to directly enhance the capabilities of LLMs for TS tasks. For example, Cao et al. (2024) captures complex interactions among trend, seasonal, and residual components to aid distribution adaptation. Chuang et al. (2024) propose a statistical prompting strategy to enhance the performance. However, these methods overlook an important step: enable LLMs to understand TS inputs first. Other methods aim to enable LLMs to understand TS data and activate their potential for TS tasks. For example,  Jin et al. (2024) reprograms input TS into text prototypes and enhances them with prompts. Sun et al. (2024) aligns TS embeddings space with LLMs embeddings space. Pan et al. (2024) proposes S2IP-LLM, a semantic space-informed prompt learning to align TS embeddings with language token embeddings. However, these methods overlook that LLMs’ inherent strength stems from their understanding of the logic and structure, rather than superficial token embedding processing. Merely aligning token embeddings without considering the coherence and consistency of the context fails to leverage this inherent advantage.
@@ 3.2 Demo: Vanilla Context-Alignment (VCA)
The most straightforward approach to utilizing pre-trained LLMs for TS tasks is to input both TS data and vanilla language prompts into the model directly. For TS forecasting tasks, a vanilla prompt like “Predict future sequences using previous data:" can be employed to guide LLMs in completing the prediction. The embeddings of the prompt can be represented as $\{{\bm{z}}_{1},{\bm{z}}_{2},\dots,{\bm{z}}_{m}\}$, and the overall input embeddings of the model can be represented as below:
@@
Due to the verbose and lack of clear structural divisions of TS embeddings $\{\bm{e}_{i}\}_{i=1}^{n}$, LLMs lack information that TS embedding $\{\bm{e}_{i}\}_{i=1}^{n}$ is an integral entity, making it difficult to analyze TS. Furthermore, form 1 lacks logical guidance. Directly concatenating the TS embeddings and language prompt embeddings loses essential contextual coherence between TS data and the prompt.
Dual-Scale Context-Alignment GNNs of VCA. Context-Alignment leverages the LLMs’ inherent strength in logic and structure to enable LLMs to contextualize and comprehend TS data. We delineate clear structures (structural alignment) and establish correct logical guidance (logical alignment) by a Dual-Scale Context-Alignment GNNs framework to achieve Context-Alignment. Firstly, using dual-scale nodes, structural alignment aggregates tokens from the same modality into one linguistic component while preserving the feature of each token. Fine-grained GNN $G_{F}$ treats each token, i.e., each element in the form 1 as a node. Coarse-grained GNN $G_{C}$ treats consecutive tokens with the same modality as a node. Specifically, $G_{C}$ employs two learnable linear layers, $f_{e}$ and $f_{z}$, to embed the TS tokens and language tokens into an $M$-dimensional space, respectively, which can be formalized as
@@
Each element in the form 2 is regarded as a node in $G_{C}$. Secondly, using directed edges, logical alignment emphasizes the correct semantic association between different components. Clearly, in the prompt “Predict future sequences using previous data:", “previous data:” refers to the lengthy TS data, and requires information from TS data. Therefore, in $G_{C}$, we construct directed edge $E:\tilde{\bm{e}}\to\tilde{\bm{z}}$ to indicate that the entire TS data serves as the upstream information source for the prompt. In $G_{F}$, we construct directed edges $\{E_{ij}:{\bm{e}}_{i}\to{\bm{z}}_{j}|i\in\{1,\dots,n\}$, $j\in\{1,\dots,m\}\}$ to convey varying information from each TS token to the prompt token. We also constrain the sum of edge weights from TS tokens to one language token to be $1$, i.e. $\sum_{i=1}^{n}w_{ij}=1$, $w_{ij}$ is the edge weight of $E_{ij}$, implicitly emphasizing that TS should be treated as a whole. In $G_{C}$, the weights of all directed edges are set to $1$. Besides, $\{w_{ij}\}_{i=1}^{n}$ are proportional to the cosine similarity between the embeddings of two nodes.
@@
Overview. After structural and logical alignment, $\hat{\bm{N}}_{F}$ maintains clear structural and coherent semantics information, helping LLMs to comprehend TS tasks and activating potential capabilities. Both $\hat{\bm{N}}_{F}$ and $\hat{\bm{N}}_{C}$ are input into pre-trained LLMs. The DSCA-GNNs can be flexibly integrated into various layers of pre-trained LLMs, and only the first time apply it need $f_{e}$ and $f_{z}$ to obtain coarse-grained GNN. The output from the $G_{F}$ branch is used to compute the MSE loss against the ground truth. In this section, we use the vanilla prompt to construct a specific DSCA-GNNs framework, we call it Vanilla Context-Alignment (VCA).
3.3 Few-Shot Prompting Based Context-Alignment (FSCA)
@@
In NLP, “few-shot prompting" refers to a small set of instances provided to the model to demonstrate a task, enabling the model to perform similar tasks effectively (Brown, 2020). Based on this, we divide the TS embeddings $\{\bm{e}_{i}\}_{i=1}^{n}$ into $N$ parts while preserving their original order. Since the subsequent TS part can be used as ground truth for predictions based on the preceding TS parts, we can construct $N-1$ prediction demonstration examples using $N$ parts of $\{\bm{e}_{i}\}_{i=1}^{n}$ and language prompt embeddings $\{{\bm{z}}_{1},{\bm{z}}_{2},\dots,{\bm{z}}_{m}\}$ same as demo in Sec. 3.2. The $j$-th part of $\{\bm{e}_{i}\}_{i=1}^{n}$ is denoted as $\{{\bm{e}}_{j,1},\dots,{\bm{e}}_{j,l_{j}}\}$. We arrange the TS-language embeddings in the format as:
@@
Dual-Scale Context-Alignment GNNs of FSCA.  Due to the verbose and lack of clear structural divisions of the format 5, LLMs struggle to understand the input. We first construct the coarse-grained GNN $G_{C}$. We introduce two learnable linear layers $f_{e}$ and $f_{z}$ to embed the TS tokens and language tokens into an $M$-dimensional space, respectively, which can be formalized as
@@
Fine-grained GNN $G_{F}$ treats each token in the form 5 as a node. The directed edge set of the $G_{F}$ can be represented as follows:
@@
The formula LABEL:edgeF means that the directed edges of $G_{F}$ are a decomposition of the directed edges in $G_{C}$. Additionally, since LLMs have strong comprehension abilities for language prompts, we can prune the directed edges in $G_{F}$, transforming form LABEL:edgeF into form LABEL:edgeF2, i.e., TS tokens are only connected to the first and last tokens of the prompt, thereby prevent overfitting. We constrain $\sum_{s=1}^{l_{j}}w^{(i)}_{j,s}=1$ for the first type edges, $\sum_{s=1}^{l_{i+1}}w^{(i)}_{i+1,s}=1$ for the second type edges, where $w^{(i)}_{j,s}$ is the edge weight of ${\bm{e}}_{j,s}\to{\bm{z}}^{(i)}_{1}$, $w^{(i)}_{i+1,s}$ is the edge weight of ${\bm{z}}^{(i)}_{m}\to{\bm{e}}_{i+1,s}$. $\{w^{(i)}_{i+1,s}\}_{s=1}^{l_{j}}$ and $\{w^{(i)}_{i+1,s}\}_{s=1}^{l_{i+1}}$ are proportional to the cosine similarity between node embeddings. Fig. 1 provides a schematic of the structure. The node embedding update formula is similar to formula 3.
@@
Learnable interactions & Overview. Similar to the demo in Sec. 3.2, we introduce an assignment matrix and a learnable weight matrix to achieve learnable interactions between the two scales. The training of FSCA can refer to Sec. 3.2 and Fig. 1.